OpenStudy (anonymous):

Hi, I need help resolving a derivative using Definition of Derivative method Hi, I need help resolving a derivative using Definition of Derivative method @Calculus1

OpenStudy (anonymous):

the exercise is: \[\sqrt{9x^{2}+16}\]

OpenStudy (anonymous):

this is what I've done: \[(\sqrt{9x ^{2}+18hx+9h ^{2}+16} - \sqrt{9x ^{2}+16} ) \div h\]

OpenStudy (anonymous):

have you tried rationalizing by taking the conugate?

OpenStudy (asnaseer):

Using hint from Outkast3r09:\[f(x)=\sqrt {9x ^{2}+16}\]Definition of derivative is\[f\prime(x)=\lim_{h \rightarrow 0}{(f(x+h)-f(x))/h}=\lim_{h \rightarrow 0}{(\sqrt {9(x+h) ^2+16}-\sqrt {9x ^{2}+16})/h}\]\[f\prime(x)=\lim_{h \rightarrow 0}{(\sqrt {9(x+h) ^2+16}-\sqrt {9x ^{2}+16})*(\sqrt {9(x+h) ^2+16}+\sqrt {9x ^{2}+16})/h(\sqrt {9(x+h) ^2+16}+\sqrt {9x ^{2}+16})}\]\[f\prime(x)=\lim_{h \rightarrow 0}{(9(x+h) ^{2}+16-9x ^{2}-16)/h(\sqrt {9(x+h) ^2+16}+\sqrt {9x ^{2}+16})}\]\[f\prime(x)=\lim_{h \rightarrow 0}{(18hx+9h ^{2})/h(\sqrt {9(x+h) ^2+16}+\sqrt {9x ^{2}+16})}\]\[f\prime(x)=\lim_{h \rightarrow 0}{(18x+9h)/(\sqrt {9(x+h) ^2+16}+\sqrt {9x ^{2}+16})}\]\[f\prime(x)=(18x)/2\sqrt {9x ^{2}+16}\]\[f\prime(x)=9x/\sqrt {9x ^{2}+16}\]